OpenStudy (anonymous):

Simplify each expression. Assume that all variables represent positive real numbers. (-x\frac{ 2}{ 3 } times x\frac{ 1}{ 2 })div (x times -x\frac{ 3}{ 2 } )

OpenStudy (anonymous):

\[(-x\frac{ 2}{ 3 } \times x\frac{ 1}{ 2 })\div (x \times -x\frac{ 3}{ 2 } )\]